OpenStudy (anonymous):

show that 2^(n-1) - 2^(n-2) = 2^(n-2) . please show me the process thank you :)

OpenStudy (anonymous):

\[2^{n-2}[2-1]=2^{n-2}\]

OpenStudy (anonymous):

factor out 2^(n-2)... \(\large 2^{n-1} - 2^{n-2} = 2^{n-2} \) \(\large 2^{n-2}(2^1 - 1) = 2^{n-2} \) \(\large 2^{n-2}(2 - 1) = 2^{n-2} \)

OpenStudy (anonymous):

Take 2^{n-2} common outside!

OpenStudy (shamim):

\[2^{(n-1)}-2^{(n-2)}\]

OpenStudy (shamim):

\[or, 2^{n}/2-2^{n}/4\]

OpenStudy (shamim):

\[or, 2^{n}(1/2-1/4)\]

OpenStudy (anonymous):

thank you so much for all the answers, understood it now ! really thank you ^^

OpenStudy (shamim):

\[2^{n}((2-1)/4\]

OpenStudy (shamim):

\[or, 2^{n}/4\]

OpenStudy (shamim):

\[or, 2^{n}/2^{2}\]

OpenStudy (shamim):

\[=2^{n-2}\]